A corporation issues a bond that generates the above cash flows. If the periods represent 6-month intervals, which of the following best describes this bond?0 1 2 3 59 60 $87.50 $87.50 $87.50 $87.50 $5,087.50 A corporation issues a bond that generates the above cash flows. If the p
Answer #1

Coupon payments are made periodically

As inferred from diagram, semi annual coupon payment = $87.50

And face value = 5,087.5 -87.5 =$5,000

Hence, semi annual coupon rate = 87.5/5000 = 1.75%

Annual coupon rate = 1.75%*2 = 3.5%

Coupon rate is always stated in annual terms

Hence, the answer is b. A 30 year bond with face value of $5,000 and coupon rate of 3.5% paid semi annually